# DHGCN: Dynamic Hop Graph Convolution Network for Self-supervised Point Cloud Learning

### 05 Jan 2024 v1

@@ 1 Introduction
2 Related Work
2.1 Self-supervised Point Cloud Learning
@@
2.2 Graph-based Learning Methods
@@ 3.1 Volumetric Partition
3.2 Part Feature Extraction
@@
and $p_{i}$ contains the indices of points in the $i$-th part as described in Section 3.1.
3.3 Hop Graph Convolution Module
@@
Since the low-level geometric distance can be represented by the degree of adjacent relationship between parts, we extract it by reconstructing the hop distance matrix of $\tilde{G}$ with the proposed self-supervised task. We denote the predicted hop distance of the $i$-th and $j$-th nodes of $\tilde{G}$ as $\tilde{D}_{ij}$ and consider predicting $\tilde{D}_{ij}$ as a classification problem of $\delta+1$ categories (including self-loops), where $\delta$ is the maximum distance as defined in Section 3.1. In this sense, we predict $\tilde{D}_{ij}$ by applying the MLP $g_{h}$ to edge features $e_{ij}^{\prime}$ as:
@@
4 Experimental Results
4.1 Experimental Setting

### 20 Jan 2024 v2

@@ Introduction
Related Work
Self-supervised Point Cloud Learning
@@
Graph-based Learning Methods
@@
Part Feature Extraction
@@
and $p_{i}$ contains the indices of points in the $i$-th part.
Hop Graph Convolution Module
@@
Since the low-level geometric distance can be represented by the degree of adjacent relationship between parts, we extract it by reconstructing the hop distance matrix of $\tilde{G}$ with the proposed self-supervised task. We denote the predicted hop distance of the $i$-th and $j$-th nodes of $\tilde{G}$ as $\tilde{D}_{ij}$ and consider predicting $\tilde{D}_{ij}$ as a classification problem of $\delta+1$ categories (including self-loops), where $\delta$ is the maximum distance. In this sense, we predict $\tilde{D}_{ij}$ by applying the MLP $g_{h}$ to edge features $e_{ij}^{\prime}$:
@@
Experimental Results
Experimental Setting